One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
Answer by ankor@dixie-net.com(22740)

One number is 12 more than another.
Let x = one number (larger) ; then (x-12) = another number(smaller)
:
The sum of the smaller number and twice the larger number is 39.
(x-12) + 2x = 39
:
Find the larger number:
x + 2x = 39 + 12
3x = 51
x = 51/3
x = 17
:
Smaller number; 17-12 = 5.
:
5 + 2(17) = 39 satisfies the equation: